### v2

@@ Abstract
In complex production lines, it is essential to have strict, fast-acting rules to determine whether the system is In Control (InC) or Out of Control (OutC). This study explores a bio-inspired method that digitally mimics ant colony behavior to classify InC/OutC states, enabling the detection of process transitions requiring maintenance. A case study on industrial potato chip frying provides the application context. During each two-minute frying cycle, sequences of eight temperature readings are collected. Each sequence is treated as a digital “ant" depositing virtual pheromones, generating a Base Score. New sequences, representing new ants, can either reinforce or weaken this score, leading to a Modified Base Score that reflects the system’s evolving condition. Signals such as extreme temperatures, large variations within a sequence, or the detection of change-points contribute to a Threat Score, which is added to the Modified Base Score. Since pheromones naturally decay over time unless reinforced, an Environmental Score is incorporated to reflect recent system dynamics, imitating real ant behavior. This score is calculated from the Modified Base Scores collected over the past hour. The resulting Total Score, obtained as the sum of the Modified Base Score, Threat Score, and Environmental Score, is used as the main indicator for real-time system classification and forecasting of transitions from InC to OutC. This ant colony optimization-inspired approach provides an adaptive and interpretable framework for process monitoring and predictive maintenance in industrial environments.
@@ 1 Introduction
In a production line, the system is expected to remain In Control (InC) under normal operating conditions. However, a wide range of factors can lead to an Out of Control (OutC) state, often requiring production interruptions for maintenance or repair. Both false discoveries (i.e., incorrectly identifying the system as OutC), and false omissions (i.e., failing to detect a true OutC state), can result in significant financial losses. Therefore, prompt and accurate classification of the system as InC or OutC, along with the ability to predict imminent transitions from InC to OutC, is of critical importance in Statistical Quality Control (SQC). To ensure reliability, such classification systems must be evaluated not only for their speed and accuracy but also in terms of their sensitivity and specificity.
@@
Classification algorithms based on ACO have been developed by Parpinelli et al. [6], Martens et al. [7], and Otero et al. [8]. Zhang et al. [9] proposed an enhanced genetic AC algorithm for bankruptcy prediction, which is relevant to our forecasting context. The quality of the classification results is evaluated using confusion and matching matrices, which capture the agreement and non-agreement between predicted classifications, i.e., Diagnosed InC (DInC) and Diagnosed OutC (DOutC), and the actual outcomes, namely InC (Negative) and OutC (Positive).
@@
  • False Positives: $\mathrm{FP=DOutC|InC}$ — diagnosing the system as out of control when it is actually in control;
@@
  • False Negatives: $\mathrm{FN=DInC|OutC}$ — diagnosing the system as in control when it is actually out of control.
@@
Mathews Correlation Coefficient $\rm{MCC=\sqrt{TPR\times TNR\times PPV\times NPV}\,-}$
@@
In Section 5, the classification performance is evaluated. This evaluation is based on false discoveries and false omissions measures derived from the confusion and matching matrices. Section 6 summarizes the main findings, with a comparison to the authors’ earlier results, which relied on less structured internal data from 2024; it also discusses the observed limitations. These reflections brings to mind Sir Ronald Fisher’s well-known aphorism at the First Session of the Indian Statistical Conference in Calcutta (1938): “To consult the statistician after an experiment is finished is often merely to ask him to conduct a post mortem examination".
@@ 2 The data
On the other hand, the rheostat regulating the Joule heating effect can sometimes introduce excessive corrections, namely by cooling the oil too much after temperature spikes or overheating after persistently low temperature sequences, such as when all readings fall below 180°. Since production supervisors are not particularly inclined to work directly with numerical data, historical practices included the use of visual indicators based on an ordinal scale inspired by Stevens [1]. Within the framework of increasing temperatures, this system used color-coded annotations to represent sequences:
@@
  • Temperatures exceeding 188° were marked red;
@@
These color classifications were grounded in the properties of the Gaussian distribution. That is, the thresholds were chosen to correspond to typical deviations from the mean in the InC state allowing for intuitive visual cues: orange indicates values within one standard deviation from the mean, red signals unusually high values (greater than two standard deviations), blue highlights consistently below-average conditions, and violet flags extremely low initial values, which is often indicative of system anomalies or start-up instability.
@@
Sequences were also annotated for extreme values, defined as maxima exceeding 195°, minima below 174°, or a range greater than 13°. These annotations are displayed alongside each sequence, as illustrated in Figure 1, which shows a typical data visualization as seen on the supervisor’s monitor screen.
@@ 3 Shortcomings of classical SQC methods
SQC is a well-established discipline within Statistics. In an ideal setting, the application of methodologies described in good textbooks, such as Montgomery [25], together with the use of robust analytical software, would allow for a rigorous and balanced diagnosis of weaknesses in factory data analysis and the formulation of appropriate and effective corrective actions. In practice, however, the realities of data collection fall far short of this ideal. As De Veaux and Hand [26] emphatically observe in the opening line of their paper’s introduction, “Bad data can ruin any analysis." They further note that “Anyone who has analyzed real data knows that the majority of their time on a data analysis project will be spent ‘cleaning’ the data before doing any analysis." We show that, in this setting, classical methodologies fail not because the methodologies themselves are flawed, but because of the presence of messy data.
Figure 2 displays the daily temperature readings (n=4,800n=4{,}800) collected on January 2, a day on which no production halt occurred. Approximately 72% of the recorded temperatures fall within the nominal $1\sigma$ limits ($176^{\degree},184^{\degree}]$), while 0.29% fall outside the nominal $3\sigma$ limits ($168^{\degree},192^{\degree}]$). These statistics reflect the observed distribution of measurements and may deviate from the theoretical expectations under a Gaussian assumption. Moreover, due to the continuous nature of the frying process, successive temperature observations exhibit temporal dependence, resulting in inherent autocorrelation (lag-1 autocorrelation = 0.634).
@@ 4 In-control/Out-of-control classification using digital pheromones
In adapting this approach to the potato chip frying process, the InC/OutC classification task was addressed. Each sequence of 8 temperature readings, captured over a 2-minute frying interval, is treated as a digital ant that deposits a trail of digital pheromones. A transition from InC to OutC is signaled as a potential vulnerability in the production line. Drawing inspiration from CVSS, a set of evaluation metrics were defined: the Base Score, Modified Base Score, Threat Score, and Environmental Score. These metrics contribute to the computation of a composite digital pheromone Total Score, which is detailed below. To clarify the structure of the scoring and classification process, Figure 4 provides a schematic overview.
@@ 4.1 Base Score
If all temperatures lie within the interval $180^{\degree},184^{\degree}]$, indicating that the process is in an InC state, then $\text{BS}=0$. Conversely, when all temperature readings exceed 192°, the multiplier $\frac{N_{184\uparrow}+N_{188\uparrow}+N_{192\uparrow}-N_{180\downarrow}}{2}$ reaches its maximum value of 12. Given that, under an InC state, the ratio $\frac{\max}{\min}\geq 1$ and is unlikely to exceed $\frac{\mu+3\sigma}{\mu-3\sigma}=1+\frac{6\sigma}{\mu-3\sigma}\approx 1.1429$, BS values greater than or approaching 12 provide a strong indication that the process might be operating in an OutC state.
@@
(1) Discard, as it likely results from delayed feedback and control during startup.
@@ 4.2 Modified Base Score
(1) Discard, as it likely results from delayed feedback and control during startup.
@@ 4.4 Environmental Score
The Environmental Score (ES) is defined as a sum of weighted averages of the MBS values from the previous hour (i.e., for the last 30 “ants"), incorporating the decay of digital pheromones over time. It is computed as
@@ 5 DInC/DOutC classification using digital pheromones
Table 7: Alarm activation
dates and times
@@
(1) Training. (2) False Discovery. (3) False Omission.
(4) Testing.
Figure 12: Production interruption between January and April
The complementary contributions of BS, MBS, ES, and TS yield the confusion matrix and related performance metrics presented in Table 8. This result is considered appropriate, given the high values of both sensitivity and specificity, and the fact that the sensitivity-specificity trade-off appropriately emphasizes sensitivity.
Table 8: Confusion matrix (March-April)
@@ 6 Conclusions
The factory’s primary objective was to predict imminent transitions from InC to OutC, and our initial intervention therefore focused on advising how better data could be obtained for this task, given the low-quality records observed in 2024 when we were first consulted. However, the data collected between January and April 2025 remained messy. In this context, our digital ant-pheromone-inspired methodology proved effective for InC/OutC classification, but it could not elicit any reliable, actionable rule for forecasting when a halt for maintenance or repair would be needed. This is a central limitation of the present study because classification performance does not automatically translate into forecasting.
@@
Operational observations further illustrate this challenge. The rheostat controlling the electric current can rapidly adjust frying temperatures; even when temperatures exceed 195°, it appears capable of returning to the target range. Yet production interruptions on January 24, February 15 (twice), April 16, and April 29 followed sequences of moderately elevated temperatures that did not trigger timely corrective adjustments. At the same time, the introduction of additional anomaly-correction devices may have been excessive. An unexpectedly large number of “ant" sequences exhibited all 8 temperature readings below the 180° threshold, producing undercooked batches and undermining crispness, de-oiling efficiency, and seasoning quality. These mixed signals, namely episodes of elevated temperatures without timely correction, alongside frequent low-temperature sequences consistent with overcorrection, are precisely the kind of confounding patterns that make forecasting rules fragile when the underlying data stream is noisy.
@@
Despite these limitations, several elements were clearly valuable in practice. The supervision team reported that the real-time temperature display, enhanced with color coding for surges (both high and low), extreme values, and change-point detections, was highly effective for visually tracking process evolution. They also considered sound and vibration alerts, activated when Base Score, Modified Base Score, Environmental Score, or Total Score indicated a possible shift towards an OutC state, appropriate and useful. Beyond immediate operational relevance, this case study demonstrates the promise of digital ants and digital pheromones, concepts from AC superorganism modeling, as ingredients in SQC classification algorithms, particularly as a way to aggregate weak local signals into a global indication of process state.
@@
It is important, however, to emphasize that this first application of digital scoring in an SQC context remains exploratory. The digital pheromone scores are informed by prior experience with markers that distinguish InC from OutC states, and this reliance on expert-informed structure is both a strength and a limitation. While the Total OutC Score generally yielded accurate classifications, the current formulation of the Threat Score did not significantly improve decision-making. This suggests that some components of the scoring architecture may require reformulation and that apparent signal may, in part, reflect artifacts of the data collection process rather than stable precursors of true OutC events.
@@
Declarations
  • Acknowledgements: The authors are grateful for the permission to use and archive the 2025 data they were asked to analyze, provided that the company’s trademark remain undisclosed to protect its reputation.
  • Funding: This work is funded by national funds through FCT — Fundação para a Ciência e a Tecnologia, I.P., under CIAC Research Unit, UID/04019/2025, DOI: https://doi.org/10.54499/UID/04019/2025, and under CEAUL Research Unit, UID/00006/2025, DOI: https://doi.org/10.54499/UID/00006/2025.
  • Conflict of interest: The authors declare no conflict of interest.
  • Author contribution: The authors contributed equally to this work.

### v3

@@ Abstract
In complex production lines, it is essential to have strict, fast-acting rules to determine whether the system is In Control (InC) or Out of Control (OutC). This study explores a bio-inspired method that digitally mimics ant colony behavior to classify InC/OutC states, enabling the detection of process transitions requiring maintenance. A case study on industrial potato chip frying provides the application context. During each two-minute frying cycle, sequences of eight temperature readings are collected. Each sequence is treated as a digital “ant” depositing virtual pheromones, generating a Base Score. New sequences, representing new ants, can either reinforce or weaken this score, leading to a Modified Base Score that reflects the system’s evolving condition. Signals such as extreme temperatures, large variations within a sequence, or the detection of change-points contribute to a Threat Score, which is added to the Modified Base Score. Since pheromones naturally decay over time unless reinforced, an Environmental Score is incorporated to reflect recent system dynamics, imitating real ant behavior. This score is calculated from the Modified Base Scores collected over the past hour. The resulting Total Score, obtained as the sum of the Modified Base Score, Threat Score, and Environmental Score, is used as the main indicator for real-time system classification and forecasting of transitions from InC to OutC. This ant colony optimization-inspired approach provides an adaptive and interpretable framework for process monitoring and predictive maintenance in industrial environments.
@@ 1 Introduction
In a production line, the system is expected to remain In Control (InC) under normal operating conditions. However, a wide range of factors can lead to an Out of Control (OutC) state, often requiring production interruptions for maintenance or repair. Both false discoveries (i.e., incorrectly identifying the system as OutC) and false omissions (i.e., failing to detect a true OutC state) can result in significant financial losses. Therefore, prompt and accurate classification of the system as InC or OutC, along with the ability to predict imminent transitions from InC to OutC, is of critical importance in Statistical Quality Control (SQC). To ensure reliability, such classification systems must be evaluated not only for their speed and accuracy but also in terms of their sensitivity and specificity.
@@
Classification algorithms based on ACO have been developed by Parpinelli et al. [23], Martens et al. [16], and Otero et al. [22]. Zhang et al. [39] proposed an enhanced genetic AC algorithm for bankruptcy prediction, which is relevant to our forecasting context. The quality of the classification results is evaluated using confusion and matching matrices, which capture the agreement and non-agreement between predicted classifications, i.e., Diagnosed InC (DInC) and Diagnosed OutC (DOutC), and the actual outcomes, namely, InC (Negative) and OutC (Positive).
@@
  • False Positives: $\mathrm{FP=DOutC|InC}$ — diagnosing the system as out-of-control when it is actually in control;
@@
  • False Negatives: $\mathrm{FN=DInC|OutC}$ — diagnosing the system as in-control when it is actually out of control.
@@
Matthews Correlation Coefficient $\rm{MCC=\sqrt{TPR\times TNR\times PPV\times NPV}}$
@@
In Section 5, the classification performance is evaluated. This evaluation is based on false discoveries and false omissions measures derived from the confusion and matching matrices. Section 6 summarizes the main findings, with a comparison to the authors’ earlier results, which relied on less structured internal data from 2024; it also discusses the observed limitations. These reflections bring to mind Sir Ronald Fisher’s well-known aphorism at the First Session of the Indian Statistical Conference in Calcutta (1938): “To consult the statistician after an experiment is finished is often merely to ask him to conduct a post-mortem examination.
@@ 2 The data
On the other hand, the rheostat regulating the Joule heating effect can sometimes introduce excessive corrections, namely, by cooling the oil too much after temperature spikes or overheating after persistently low temperature sequences, such as when all readings fall below 180°. Since production supervisors are not particularly inclined to work directly with numerical data, historical practices included the use of visual indicators based on an ordinal scale inspired by Stevens [29]. Within the framework of increasing temperatures, this system used color-coded annotations to represent sequences:
@@
  • Temperatures exceeding 188° were marked in red;
@@
These color classifications were grounded in the properties of the Gaussian distribution. That is, the thresholds were chosen to correspond to typical deviations from the mean in the InC state, allowing for intuitive visual cues: orange indicates values within one standard deviation from the mean, red signals unusually high values (greater than two standard deviations), blue highlights consistently below-average conditions, and violet flags extremely low initial values, which is often indicative of system anomalies or start-up instability.
@@
Sequences were also annotated for extreme values, defined as maxima exceeding 195°, minima below 174°, or a range greater than 13°. These annotations are displayed alongside each sequence, as illustrated in Figure 1, which shows a typical data visualization as seen on the supervisors’ monitor screens.
@@ 3 Shortcomings of classical SQC methods
SQC is a well-established discipline within Statistics. In an ideal setting, the application of methodologies described in good textbooks, such as Montgomery [19], together with the use of robust analytical software, would allow for a rigorous and balanced diagnosis of weaknesses in factory data analysis and the formulation of appropriate and effective corrective actions. In practice, however, the realities of data collection fall far short of this ideal. As De Veaux and Hand [5] emphatically observe in the opening line of their paper’s introduction, “Bad data can ruin any analysis.” They further note that “Anyone who has analyzed real data knows that the majority of their time on a data analysis project will be spent ‘cleaning’ the data before doing any analysis.” We show that, in this setting, classical methodologies fail not because the methodologies themselves are flawed, but because of the presence of messy data.
Figure 2 displays the daily temperature readings (n=4,800n=4{,}800) collected on January 2, a day on which no production halt occurred. Approximately 72% of the recorded temperatures fall within the nominal $1\sigma$ limits ($176\degree,184\degree]$), while 0.29% fall outside the nominal $3\sigma$ limits ($168\degree,192\degree]$). These statistics reflect the observed distribution of measurements and may deviate from the theoretical expectations under a Gaussian assumption. Moreover, due to the continuous nature of the frying process, successive temperature observations exhibit temporal dependence, resulting in inherent autocorrelation (lag-1 autocorrelation =0.634=0.634).
@@ 4 In-control/out-of-control classification using digital pheromones
In adapting this approach to the potato chip frying process, the InC/OutC classification task was addressed. Each sequence of 8 temperature readings, captured over a 2-minute frying interval, is treated as a digital ant that deposits a trail of digital pheromones. A transition from InC to OutC is signaled as a potential vulnerability in the production line. Drawing inspiration from CVSS, a set of evaluation metrics was defined: the Base Score, Modified Base Score, Threat Score, and Environmental Score. These metrics contribute to the computation of a composite digital pheromone Total Score, which is detailed below. To clarify the structure of the scoring and classification process, Figure 4 provides a schematic overview.
@@ 4.1 Base score
If all temperatures lie within the interval $180\degree,184\degree]$, indicating that the process is in an InC state, then $\text{BS}=0$. Conversely, when all temperature readings exceed 192°, the multiplier $\frac{N_{184\uparrow}+N_{188\uparrow}+N_{192\uparrow}-N_{180\downarrow}}{2}$ reaches its maximum value of 12. Given that, under an InC state, the ratio $\frac{\max}{\min}\geq 1$ and is unlikely to exceed $\frac{\mu+3\sigma}{\mu-3\sigma}=1+\frac{6\sigma}{\mu-3\sigma}\approx 1.1429$, BS values greater than or approaching 12 provide a strong indication that the process might be operating in an OutC state.
@@
  • (1)
    Discard, as it likely results from delayed feedback and control during startup.
@@ 4.2 Modified base score
  • (1)
    Discard, as it likely results from delayed feedback and control during startup.
@@ 4.4 Environmental score
The Environmental Score (ES) is defined as a sum of weighted averages of the MBS values from the previous hour (i.e., for the last 30 “ants”), incorporating the decay of digital pheromones over time. It is computed as
@@ 5 DInC/DOutC classification using digital pheromones
Table 7: Alarm activation dates and times
@@
  • (1)
    Training.
  • (2)
    False Discovery.
  • (3)
    False Omission.
  • (4)
    Testing.
Figure 12: Production interruptions between January and April
The complementary contributions of BS, MBS, ES, and TS yield the confusion matrix and related performance metrics presented in Table 8. This result is considered acceptable, given the high values of both sensitivity and specificity, and the fact that the trade-off between them appropriately emphasizes sensitivity.
Table 8: Confusion matrix (March–April)
@@ 6 Conclusions
The factory’s primary objective was to predict imminent transitions from InC to OutC, and our initial intervention, therefore, focused on advising how better data could be obtained for this task, given the low-quality records observed in 2024 when we were first consulted. However, the data collected between January and April 2025 remained messy. In this context, our digital ant-pheromone-inspired methodology proved effective for InC/OutC classification, but it could not elicit any reliable, actionable rule for forecasting when a halt for maintenance or repair would be needed. This is a central limitation of the present study because classification performance does not automatically translate into forecasting.
@@
Operational observations further illustrate this challenge. The rheostat controlling the electric current can rapidly adjust frying temperatures; even when temperatures exceed 195°, it appears capable of returning to the target range. Yet production interruptions on January 24, February 15 (twice), April 16, and April 29 followed sequences of moderately elevated temperatures that did not trigger timely corrective adjustments. At the same time, the introduction of additional anomaly-correction devices may have been excessive. An unexpectedly large number of “ant” sequences exhibited all 8 temperature readings below the 180° threshold, producing undercooked batches and undermining crispness, de-oiling efficiency, and seasoning quality. These mixed signals, namely, episodes of elevated temperatures without timely correction, alongside frequent low-temperature sequences consistent with overcorrection, are precisely the kind of confounding patterns that make forecasting rules fragile when the underlying data stream is noisy.
@@
Despite these limitations, several elements were clearly valuable in practice. The supervision team reported that the real-time temperature display, enhanced with color coding for surges (both high and low), extreme values, and change-point detections, was highly effective for visually tracking process evolution. They also considered the sound and vibration alerts, activated when Base Score, Modified Base Score, Environmental Score, or Total Score indicated a possible shift towards an OutC state, to be appropriate and useful. Beyond immediate operational relevance, this case study demonstrates the promise of digital ants and digital pheromones, concepts from AC superorganism modeling, as ingredients in SQC classification algorithms, particularly, as a way to aggregate weak local signals into a global indication of process state.
@@
It is important, however, to emphasize that this first application of digital scoring in an SQC context remains exploratory. The digital pheromone scores are informed by prior experience with markers that distinguish InC from OutC states, and this reliance on expert-informed structure is both a strength and a limitation. While the Total OutC Score generally yielded accurate classifications, the current formulation of the Threat Score did not significantly improve decision-making. This suggests that some components of the scoring architecture may require reformulation and that the apparent signal may, in part, reflect artifacts of the data collection process rather than stable precursors of true OutC events.